OpenStudy (anonymous):

7/20 of the animals on a farm are sheep and 2/5 are cattle. a . What percentage of the animals are sheep ? b. What percentage of the animals are cattle ? c. If there are 480 animals how many are sheep or cattle ?

OpenStudy (anonymous):

A. \[(7\div20)*100\]

OpenStudy (anonymous):

B.\[\frac{ 2 }{ 5 }\times100\]

OpenStudy (anonymous):

C. Once you get the percent in A and B, divide it by 100 and then multiply the quotient with 480.

OpenStudy (anonymous):

Okay , thank you so much for your help :)

OpenStudy (anonymous):

I'm not sure with C.

OpenStudy (anonymous):

Its alright , dont worry :) You have helped me a lot anyway

OpenStudy (anonymous):

C. Or you can make it as proportion. :) Like: \[\frac{ 7}{ 20 }\times \frac{ x }{ 480 }\]

OpenStudy (anonymous):

Okay :D

OpenStudy (anonymous):

Either way, the answer is the same. :)

OpenStudy (anonymous):

Ok :)